suppose that the travel time from your home to your office is normally distributed with mean 40 min and standard deviation 7 min. if you want to be 90% certain that you will not be late for an appointment at 1, what is the latest time that you should leave home?
Answer #1

Given that,

mean = = 40

standard deviation = = 7

Using standard normal table,

P(Z < z) = 90%

= P(Z < z) = 0.90

= P(Z < 1.28) = 0.90

z = 1.28   

Using z-score formula  

x= z * +

x= 1.28 *7+40

x= 48.96

The latest time that you should leave home is 48.96